Solving Exponential Equations Practice:

  1. 8x-3 =85
  2. 103x+2 = 1011
  3. 46x = 169
1
Expert's answer
2021-05-17T16:18:02-0400

1.


"8^{x-3}=8^5"

Exponential equation function is monotone


"x-3=5"

"x=8"

2.


"10^{3x+2}=10^{11}"

Exponential equation function is monotone


"3x+2=11"

"3x=9"

"x=3"

3.


"4^{6x}=16^9"

"4^{6x}=(4^2)^9"

"4^{6x}=4^{2\cdot9}"

"4^{6x}=4^{18}"

Exponential equation function is monotone


"6x=18"

"x=3"